OpenStudy (mackenzie2013):

What is the value of the 11th term in the sequence -3, -6, -12, -24, ...? -6,144 -118,098 -3,072 -354,294

OpenStudy (the_fizicx99):

You can see that its being multiplied by 2 so the ratio is 2. With the given information just plug in into the formula: a_n = a1 • r^n-1 a1 = -3 r = 2 Plug in 11 for n, a_11 = -3 • 2^11-1 a_11 = -3• 2^10 Can you continue?

OpenStudy (mackenzie2013):

-3 x 20?

OpenStudy (mackenzie2013):

No no

OpenStudy (mackenzie2013):

2 10 times

OpenStudy (the_fizicx99):

No, 2^10 = 2 * 2 * 2 * 2 * 2 * 2 * 2 * 2 * 2 * 2 2^10 isn't 2 * 10

OpenStudy (mackenzie2013):

3,072

OpenStudy (the_fizicx99):

-3,072 *

OpenStudy (the_fizicx99):

-3 * 1,024 = -3,072

OpenStudy (mackenzie2013):

ahh sorry forgot the negative

OpenStudy (the_fizicx99):

yup, good job ;p